This directory also contains a configuration file whizzy.el which will be automatically loaded. However, local configuration, i.e. the above line, takes priority declarations in file whizzy.el. To see this, you may the line above (or simply add a white space in front of it), and observe that the declaration in whizzy.el will indeed be taken into account.

This directory also contains a file whizzy.sty, which is automatically loaded

before executing \begin{document}.

This directory also contains a configuration file whizzy.sh. This file will be loaded every time WhizzyTeX is run. To see this, look at the initialization log file (you may use the menu entry View log).

Usually, you do not need any of these configuration files.

You may try the menu entry Customize slice and enter \Large. The latex command \Large will then be inserted automatically at the beginning of every slice, and in this case let your document be typeset in bigger font. To cancel it, just let the slice customozation be the empty string.

1 First section

Some text in the section.

1.1 Sub

Some text in the subsection.

1.1.1 SubSub

Some text in the sub-subsection.

2 Foo

2.1 Bar

Par And a paragraph.